OpenStudy (anonymous):

If you have knowledge of differential equation please help me understand this one; I was given this: d^2y/dx^2 + 2dy/dx = 5 and without showing how the complementary function that they got was: CF: y = Acos(sqrt(2)*x) +Bsin(sqrt(2)*x) Please enlighten me if you know how this happened

OpenStudy (sirm3d):

the complementary function is obtained from d^2y/dx^2 + 2y = 0, NOT d^2y/dx^2 +2dy/dx = 0

OpenStudy (anonymous):

can you please show me the steps taken..and take into consideration the "5" that's been equated by the eqn

OpenStudy (sirm3d):

assuming that the DE is \[\frac{ d^2y }{ dx^2 }+2\frac{ dy }{ dx }=5\]

OpenStudy (sirm3d):

the complementary function is \[y _{c}=c _{1}e^{-2x}+c_{2}\]

OpenStudy (anonymous):

that's what I got too, guess I'll have to check in with my teacher

OpenStudy (anonymous):

Thanks by the way